The place value of a digit at the hundredths place is 1/10 times the same digit at the tenths place.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False

Let ‘a’ be the same digit at tensand hundreds place in a number.
Place value of digit at tens place = 10 × a = 10a
Place value of digit at hundreds place = 100 × a = 100a
Hence, The place value of a digit at the hundreds place is 10 times the same digit at the tens place.


Thanks